OpenStudy (anonymous):

Solve exactly log (base 3)(2x-3) = 0 x= ?

OpenStudy (anonymous):

Good to see you again jahtoday

OpenStudy (anonymous):

I saw your post on my other question and felt very stupid lol

OpenStudy (anonymous):

\[log_b(a) = 0 \implies a = b^0 = 1\]

OpenStudy (anonymous):

So how does that apply to your problem here?

OpenStudy (anonymous):

2x-3=3^0 ?

OpenStudy (anonymous):

lol sry just trying to point in the right direction and yes thatsright

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

And 3^0 = 1

OpenStudy (anonymous):

so 2x-3=1 then 2x=4 so x=2?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Nicely done.

OpenStudy (anonymous):

awesome thank you both